Suppose that the demand curve is P = 75 - 4.0 What is the price elasticity of demand when the price is $25?
Answer #1

Answer

The formula is:

E = ele P

converting the inverse demand function to normal form

P=75-4Q

4Q=75-P

Q=18.75-0.25P

P=25

Q=18.75-0.25*25

Q=12.5

= -0.25 E = -0.25 * 25 -=-0.50

the elasticity of demand is -0.5